HIGH-RESOLUTION SPECTROSCOPY AND PRELIMINARY GLOBAL ANALYSIS OF C–H STRETCH-ING VIBRATIONS OF C_{2}H_{4} IN THE 3000 AND 6000 CM^{−1} REGIONS

M. A. LORONO GONZALEZ, Department of Chemistry, Universidad de Oriente, Cumaná 6101, Estado Sucre, Venezuela; \underline{V. BOUDON}, M. LOËTE, Laboratoire Interdisciplinaire Carnot de Bourgogne, UMR 5209 CNRS-Université de Bourgogne, 9. Av. A. Savary, BP 47870, F-21078 Dijon Cedex, France; M. ROTGER, M.-T. BOURGEOIS, Groupe de Spectrométrie Moléculaire et Atmosphérique, CNRS UMR 6089, Moulin de la Housse, BP 1039, Cases 16-17, F-51687 Reims Cedex 2, France; \underline{K. DIDRICHE}, M. HERMAN, Laboratoire de Chimie quantique et Photophysique, CP160/09, Faculté des Sciences, Université Libre de Bruxelles, 50 ave. Roosevelt, B-1050, Brussels, Belgium; V. A. KAPITANOV, Yu. N. PONOMAREV, A. A. SOLODOV, A. M. SOLODOV, T. M. PETROVA, V.E. Zuev Institute of Atmospheric Optics SB RAS,1, Zuev Square, Tomsk, 634921, Russia.

Ethylene is a naturally occurring compound in ambient air that affects atmospheric chemistry and global climate. The C_{2}H_{4} spectrum is available in databases only for the 1000 and 3000 cm^{−1} ranges. In this work^{a}, the ethylene absorption spectrum was measured in the 6030-6250 cm^{−1} range with the use of a high resolution Bruker IFS 125HR Fourierspectrometer and a two-channel opto-acoustic spectrometer with a diode laser. As a secondary standard of wavelengths, the methane absorption spectrum was used in both cases. A preliminary analysis was realized thanks to the tensorial formalism developed by the Dijon group that is implemented in the XTDS software package^{b}. We considered the two combination bands ν_{5} + ν_{9} and ν_{5} + ν_{11} as an interacting dyad. Parameters for the ν_{9}/ν_{11} dyad were fitted simultaneously from a re-analysis of previously recorded supersonic expansion jet FTIR data, while parameters for the v_{5} = 1 Raman level were taken from literature. More than 600 lines could be assigned in the 6030-6250 cm^{−1} region (and also 682 in the 2950–3150 cm^{−1} region) and effective Hamiltonian parameters were fitted, including Coriolis interaction parameters. The dyad features are globally quite well reproduced, even if there are still problems at high J values.

aM. A. Loroñno Gonzalez et al., J. Quant. Spectrosc. Radiat. Transfer, 111, 2265–2278 (2010).

bCh. Wenger, V. Boudon, M. Rotger, J.-P. Champion and M. Sanzharov, J. Mol. Spectrosc., 251, 102–113 (2008).